CMLA operational status:
The station is back up after the station operator replaced the faulty Newmar.
The station is down. The station operator has reported a faulty Newmar power supply when she replaced the station’s backup batteries.
IDA engineers visited the station and installed a new STS-6 on day 2019,158-11:35 (June 7).
IDA engineers visited the station to replace the KS54000 seismometer with a new STS-6. Upon installation the STS-6 sensor failed and is being sent back to the manufacturer for repair. There are no BH*00,LH*00,or VH*00 channels until the sensor is replaced.
IDA engineers visited the station and replaced the CMG-3T seismometer with an STS-5 on day 2017,175 (June 24). An infrasound sensor, weather staton and new station computers were also installed at the station.
The vertical channels from the CMG-3T seismometer were non-seismic from day 2016,238 to 2016,240. They were restored after the station operator cycled power to the sensor. However, data quality is poor after the masses for the sensor were re-centered. The station operator has been requested to cycle power to the sensor again.
An IDA engineer is onsite to upgrade the DC power system. A Setra microbarometer was installed on day 2016,061-20:34.
An IDA engineer visited the station and replaced the faulty CMG-3T seismometer. All channels from the CMG-3T seismometer have been restored.
All channels from the CMG-3T seismometer remain non-seismic due to a faulty sensor. A replacement will be needed.
The station is back on-line after an IDA engineer visited the station and restored data acquisition. However, all channels from the CMG-3T seismometer are non-seismic. An IDA engineer is onsite.
An IDA engineer upgraded the station to a Q330HR system. The station remains down due to communication problems between the digitizers and the ISP.
An IDA engineer is on-site and is upgrading the station to a Q330HR system.
The station operator reports replacing the digitizer with one shipped from UCSD. However, the station remains down due to failure of the ISP.
The station is down as of about day 2009,361 due to problems with the digitizer that are yet to be determined.
The station is back online as of about day 2008,361 after the station operator replaced the DAS with one shipped from UCSD.
The station is down due to problems with the DAS that are yet to be determined.
The station is back online as of about day 2008,286.
The station is down as of about day 2008,283 due to problems with the DAS that are yet to be determined.
The BH1-10,LH1-10,VH1-10,and HH1-10 channels from the CMG-3T seismometer had excessive spiking and were non-seismic from about day 2008,245-1453 to 2008,254-1030. They were restored after the station operator recentered the CMG-3T seismometer.
The station is back online as of about day 2008,237 after the station operator replaced the DAS.
The station is down as of about day 2008,195 due to problems with the DAS that are yet to be determined.
An IDA engineer visited the station and installed a new CMG-3T seismometer and restored the short period channels as of about day 2008,117.
The station is back online after the station operator restored the telemetry link to the station.
The station is down as of about day 2007,313, following a lightning storm that damaged equipment at the station. A visit to the station by an IDA engineer is being planned.
The station is back online as of about day 105-10:15 after the station operator replaced the DAS, restored power to the station and restored the telemetry connection to the station. The short period channels from the CMG-3T seismometer are out due to a faulty sensor.
IDA station CMLA is down following a lightning storm that knocked out power to the wellhead (day 279). IDA engineering is working with the station operator to trace the source of the problem.
The station is up after an IDA engineer upgraded the station to a MK8 system. New equipment including a CMG-3T seismometer, strong motion sensor, ISP, GPS clock, power supplies, and a Paroscientific microbarograph have been installed.
The station is down for reasons yet to be determined. Telemetry to the station is up, however data from all channels are nonseismic.
The station is up. However, broadband channels for the KS-54000 seismometer, and short period channels for the STS-2 seismometer are out due to a faulty circuit board in the DAS.
The station is down for reasons yet to be determined.
Station operator reports receiving equipment including GPS clock,routers,and computers (04/07). Station operator reports replacing the station processor (04/29). Telemetry to the station restored 05/18 after the station operator installed new routers.
Station operator reports problem with the GPS clock (Dec. 8).
Station is operational. However, only one board in DAS is working correctly. Short period channels from the STS-2 sensor are also out.
BBZ channels are functional again after DAS was rebooted remotely to correct a large offset on a RT303 board.
The vertical channels, BBZ, are down.
Station is up after the station operator cycled power to data acquisition system.
Operator reports no data from DAS. An additional board failure is suspected.
The station is up again after the station operator removed a faulty RT303 card in the DAS.
The station is down as of 2004,077 (Mar 17) due to problems with the DAS.
Telemetry down as of 2004:038-00:04 (Feb 7) due to problem with station modems. New modems shipped 2004.03.01.
Station is up again as of 2004:020-18:18 (Jan 20) after an IDA engineer replaced the power supply, a faulty RT303 card in the DAS, the DAT drive, and station modem.
Station is down. Station operator reports replacing a surge protector which was found damaged at the phone line, after which he cycled power. However, the ARS still indicates there is no data from DAS (9/24).
Telemetry down as of 2003:260-18:17 (Sept.17).
Data from the KS54000 seismometer North-South channels restored on 2003.07.12 after station operator installed new RT303D board for the digitizer.
No NRTS data from 2003:142-23:07:02 (May 22,2003) to 2003:147-12:30:34 (May 27,2003). BHZ channel out from 2003:147-12:30:34 to 2003:150-23:29 (May 27,2003 to May 30,2003). After the CMLA digitizer was rebooted at 2003:150-23:29:12, the BHZ channel returned to normal.
SP* data restored after a cable to the digitizer was reconnected.
Station operator installed RT303 cards and new modems. Telemetry restored. Data restored from the KS54000 seismometer Vertical channels on 2003.04.10 after station operator swapped cards from the channel 2 (bhe) to the channel 0 (bhz) slots. Data from the KS54000 seismometer East-West channels restored on 2003.04.30 after the station operator installed replacement RT303 card. Currently investigating problems on the North-South channel. No SP* data.
New DAT drive shipped 2002.12.19 arrived.
New EPROMs installed to correct DSP problem.
The DAS RT303 card digitizing data from the KS54000 north sensor is behaving anomalously.
Station is operating normally.
Telemetry connections to the station now made via local ISP.